# Toward Black–Scholes for Prediction Markets: A Unified Kernel and Market-Maker’s Handbook

### v1

@@ 3.3 Multi-Event Dependence: Diffusive Correlation and Co-Jumps
where $S^{\prime}_{k}=S^{\prime}(x_{t}^{k})$. The first term is the diffusive covariation; the second aggregates common (co-)jumps. Empirically, co-jumps can be detected and tested using high-frequency methods (JacodTodorov2009, ).
@@ 3.4 Prototype Derivatives (Belief–Variance, Correlation, Corridor, and First-Passage Notes)
directly paralleling classical variance swap theory in price models (Demeterfi1999, ; CarrLee2009, ; BroadieJain2008, ).
@@
where the prefactor $\big(p(1-p)\big)^{2}$ arises from $S^{\prime}(x)^{2}$ and the second term captures jump contributions to the quadratic variation of $p$ (ContTankov2004, ; CarrLee2004, ).
@@
Using (4), a short-maturity fair covariance strike integrates the instantaneous covariance; dividing by marginal variances yields a correlation strike. These instruments let market makers neutralize cross-event exposure from both diffusive correlation and co-jumps (JacodTodorov2009, ; CarrLee2009, ).
@@
A corridor contract accrues realized variance only while $p\in[a,b]$ (a “swing zone” away from the $0/1$ boundaries). Pricing proceeds either via a weighted-variance replication (when available) or by solving the PIDE with state-dependent accrual; see corridor-variance analogs in equity for guidance (Lee2008Corridor, ; Burgard2017Corridor, ).
@@
Pay a fixed amount if $p$ first hits level $h\in(0,1)$ before $T$ (or logical AND/OR across events). With $h$ mapped to $x_{h}=\operatorname{logit}(h)$, valuation uses (7) with absorbing boundary at $x=x_{h}$ and appropriate terminal/boundary conditions. Jump terms materially affect first-passage probabilities (up-crossings can occur by jump), a standard consideration in jump–diffusion settings (ContTankov2004, ; Applebaum2009, ).
@@ 3.5 General Pricing via PIDE and Numerical Treatment
with $\mu$ given by (3). Basket and multi-event claims add diffusion cross-derivatives and a multivariate jump integral with co-jump measure $\nu_{t}(d\mathbf{z})$. Under standard growth and regularity conditions, (7) is the infinitesimal generator equation for (1) and can be solved by finite-difference with fast convolution for the jump integral, Fourier methods when coefficients are constant/affine, or Monte Carlo with variance reduction (ContTankov2004, ; Applebaum2009, ).
@@
Mapping mid/bid–ask/trade streams to a belief–volatility surface requires filtering the latent $x_{t}$ from microstructure noise (e.g., state-space/Kalman variants) and estimating jump activity and co-jumps; the microstructure and high-frequency literature provides standard tools (Hasbrouck1991, ; HasbrouckBook2007, ; AMZ2005, ; JacodTodorov2009, ). In our setting, these methods are applied to log-odds increments and to cross-event panels.

### v2

@@ 3.3 Multi-Event Dependence: Diffusive Correlation and Co-Jumps
where $S^{\prime}_{k}=S^{\prime}(x_{t}^{k})$. The first term is the diffusive covariation; the second aggregates common (co-)jumps. Empirically, co-jumps can be detected and tested using high-frequency methods (Jacod and Todorov, 2009).
@@ 3.4 Prototype Derivatives (Belief–Variance, Correlation, Corridor, and First-Passage Notes)
directly paralleling classical variance swap theory in price models (Demeterfi et al., 1999; Carr and Lee, 2009; Broadie and Jain, 2008a).
@@
where the prefactor $\big(p(1-p)\big)^{2}$ arises from $S^{\prime}(x)^{2}$ and the second term captures jump contributions to the quadratic variation of $p$ (Cont and Tankov, 2004; Carr and Lee, 2004).
@@
Using (4), a short-maturity fair covariance strike integrates the instantaneous covariance; dividing by marginal variances yields a correlation strike. These instruments let market makers neutralize cross-event exposure from both diffusive correlation and co-jumps (Jacod and Todorov, 2009; Carr and Lee, 2009).
@@
A corridor contract accrues realized variance only while $p\in[a,b]$ (a “swing zone” away from the $0/1$ boundaries). Pricing proceeds either via a weighted-variance replication (when available) or by solving the PIDE with state-dependent accrual; see corridor-variance analogs in equity for guidance (Lee, 2008; Burgard, 2017).
@@
Pay a fixed amount if $p$ first hits level $h\in(0,1)$ before $T$ (or logical AND/OR across events). With $h$ mapped to $x_{h}=\operatorname{logit}(h)$, valuation uses (7) with absorbing boundary at $x=x_{h}$ and appropriate terminal/boundary conditions. Jump terms materially affect first-passage probabilities (up-crossings can occur by jump), a standard consideration in jump–diffusion settings (Cont and Tankov, 2004; Applebaum, 2009).
@@ 3.5 General Pricing via PIDE and Numerical Treatment
with $\mu$ given by (3). Basket and multi-event claims add diffusion cross-derivatives and a multivariate jump integral with co-jump measure $\nu_{t}(d\mathbf{z})$. Under standard growth and regularity conditions, (7) is the infinitesimal generator equation for (1) and can be solved by finite-difference with fast convolution for the jump integral, Fourier methods when coefficients are constant/affine, or Monte Carlo with variance reduction (Cont and Tankov, 2004; Applebaum, 2009).
@@
Mapping mid/bid–ask/trade streams to a belief–volatility surface requires filtering the latent $x_{t}$ from microstructure noise (e.g., state-space/Kalman variants) and estimating jump activity and co-jumps; the microstructure and high-frequency literature provides standard tools (Hasbrouck, 1991, 2007; Aït-Sahalia et al., 2005; Jacod and Todorov, 2009). In our setting, these methods are applied to log-odds increments and to cross-event panels.